a standard deck of cards contains 52 cards. One card is selected from the deck.

​(a)

Compute the probability of randomly selecting a seven or three.

​(b)

Compute the probability of randomly selecting a seven or three or jack.

​(c)

Compute the probability of randomly selecting a nine or diamond




1
Expert's answer
2021-02-24T07:38:38-0500

(a) P(seven or three) = P(seven) + P(three)

"= \frac{4}{52} + \frac{4}{52} \\

= \frac{8}{52} \\

= \frac{2}{13} \\

= 0.154"

(b) P(seven or three or jack) = P(seven) + P(three) + P(jack)

"= \frac{4}{52} + \frac{4}{52} + \frac{4}{52} \\

= \frac{12}{52} \\

= \frac{3}{13} \\

= 0.230"

(c) P(nine or diamond) = P(nine) + P(diamond) – P(nine "\cap" diamond)

"= \frac{4}{52} + \frac{13}{52} - \frac{1}{52} \\

= \frac{4}{13} \\

= 0.307"